OpenStudy (anonymous):

What is the equation of the line that passes through the point (4, 1) and has a slope of 5? A. y = 5x - 3 B. y = 5x - 1 C. y = 5x + 3 D. y = 5x - 19

hero (hero):

substitute x = 4 and y = 1 into each equation. Once you get to a correct solution, then you have found the correct equation.

OpenStudy (anonymous):

so it would be D?

hero (hero):

Yes

OpenStudy (anonymous):

what about this problem: What is the equation of the line that passes through the points (-1, 3) and (1, 11)? A. 4x - y = 1 B. 4x + y = 1 C. x - 4y = -13 D. 4x - y = -7

hero (hero):

Do the same thing, except substitute BOTH points in for each. It won't take as long as you think, however, you need to make sure that BOTH points work for one equation. When you find the equation where both points work, then you have found the correct equation.

hero (hero):

When you substitute the points, if one of them fails, then the whole equation automatically fails.

OpenStudy (anonymous):

So its D?

hero (hero):

Yes

OpenStudy (anonymous):

is this correct? Choose the correct description of the graph of the compound inequality: x - 7 > -14 and 5x <= 25 A. A number line with an open circle on -7, shading to the left, and a closed circle on 5, shading to the right. B. A number line with a closed circle on -7, shading to the left, and an open circle on 5, shading to the right. C. A number line with a closed circle on -7, an open circle on 5, and shading in between. D. A number line with an open circle on -7, a closed circle on 5, and shading in between. I think its A?

hero (hero):

Actually, its D again

hero (hero):

Sorry it took so long. I was on the phone with someone. I don't know how to multi-task

OpenStudy (anonymous):

thats perfectly fine, im in no hurry